OpenStudy (anonymous):

write the equation of a line with the given information: (4,8) and (16,32)

OpenStudy (shadowfiend):

Do you know how to calculate the slope from that info?

OpenStudy (anonymous):

use the slope formula m=y2-y1 --------- < divided by x2-x1 (4,8) (16,38) x1,y1 x2, y2 plug them in to find your slope. m=38-8 ----- < divided by m=3.75 or 15/4 16-8 so now you use y=mx+b - y=yalue, m=slope, x=x value & b=y-intercept one of your y values is 8 [you can use 38 if you want] 8=3.75(4)+b 8=15+b b=-7 your final formula is y=15/4x-7